# Asymmetric Graph Error Control with Low Complexity in Causal Bandits

### v1

@@ IV Intervention Selection under Uncertainty
$\Delta\bm{B}_{\bm{a}}^{t}\doteq\hat{\bm{B}}_{\bm{a}}^{t}-\bm{B}_{\bm{a}},\quad\Delta\bm{\mu}_{\bm{a}}^{t}\doteq\hat{\bm{\mu}}_{\bm{a}}^{t}-\bm{\mu}_{\bm{a}},$ (29)
where $\hat{\bm{\mu}}_{\bm{a}}^{t}$ represents the estimated mean of $X$. As analyzed in Section III-C, the weight errors could have non-zero means, due to missing edges (false negatives). With the centered error matrix defined as
@@
$\left|\left[\Delta\bm{\mu}_{\bm{a}}^{t}\right]_{N}\right|$ $=\left|\bigl{[}(\bm{I}-\hat{\bm{B}}_{\bm{a}}^{t})^{-1}\bigr{]}_{N}^{T}(\Delta\bm{B_{a}}^{t})^{T}\bm{\mu}_{\bm{a}}\right|$ (31)
$\leq\left|\bigl{[}(\bm{I}-\hat{\bm{B}}_{\bm{a}}^{t})^{-1}\bigr{]}_{N}^{T}\mathbb{E}\bigl{[}\Delta\bm{B}_{\bm{a}}^{t}\big{|}\bm{X}^{1:t}(\bm{a})\bigr{]}^{T}\bm{\mu}_{\bm{a}}\right|$
$\hskip 40.0pt+\left|\bigl{[}(\bm{I}-\hat{\bm{B}}_{\bm{a}}^{t})^{-1}\bigr{]}_{N}^{T}\bigl{(}\Delta\tilde{\bm{B}}_{\bm{a}}^{t}\bigr{)}^{T}\bm{\mu}_{\bm{a}}\right|.$ (32)
Recall that we provides an expression for the non-zero elements of $\mathbb{E}\bigl{[}\Delta\bm{B}_{\bm{a}}^{t}\big{|}\bm{X}^{1:t}(\bm{a})\bigr{]}$ in (28), which depends on the correlation between correctly declared parents and mistakenly rejected parents. However, since we are not able to know whether and where causal edges are rejected by mistake, the uncertainty bound cannot be computed at run-time.
@@
$\mathbb{P}\biggl{\{}\sigma_{\max}(\Delta\bm{B}_{\bm{a}}^{t})\geq 2\left(N^{2}+2N\right)^{1/4}\cdot\\ \sqrt{\ln\Bigl{(}\frac{2N}{\delta}\Bigr{)}\sum_{i=1}^{N}\lambda_{\max}(\bm{\Phi}_{\bm{a}}^{t}(i))}\biggr{\}}\leq\delta,$ (34)
@@
The proof relies on on a key dilation transformation, Lemma 1 and the matrix Bernstein inequality [32]. Based on Lemma 2, the following theorem provides a concentration inequality for the error of the estimated reward.
@@
Under the assumption MMSE estimation and no false negative errors, the following inequality holds:
@@
$U(\bm{X}^{1:t},\bm{a},\delta)=2\left(N^{2}+2N\right)^{1/4}\left\lVert\bigl{[}(\bm{I}-\hat{\bm{B}}_{\bm{a}}^{t})^{-1}\bigr{]}_{N}\right\rVert_{2}\cdot\\ \left\lVert\bm{\mu}_{\bm{a}}\right\rVert_{2}\cdot\sqrt{\ln\Bigl{(}\frac{2N}{\delta}\Bigr{)}\sum_{i=1}^{N}\lambda_{\max}(\bm{\Phi}_{\bm{a}}^{t}(i))}.$ (36)
@@
The proof is based on the Woodbury matrix identity [33] and Lemma 2. Since the uncertainty of reward estimation cannot be measured in the presence of FN errors, we employ (35) as an approximation for the general case. The bound on uncertainty enables us to balance between exploration and exploitation, so that the cumulative regret is minimized in the long run. In the presence of FN errors, although the uncertainty bound may not hold exactly, the observed low cumulative regret in Section VI-B suggests that the uncertainty is effectively quantified.
@@
$\bm{a}^{t+1}=\operatorname*{arg\,max}_{\bm{a}}\left\{\bigl{[}(\bm{I}-\hat{\bm{B}}_{\bm{a}}^{t})^{-1}\bigr{]}^{T}_{N}\cdot\bm{\nu}+\alpha\cdot U(\bm{X}^{1:t},\bm{a},\delta)\right\},$ (37)
@@ V Causal Sub-graph Change Detection
So far, we have investigated the causal bandit problem under the assumption that the underlying causal structure is invariant. However, in real-world scenarios it could be non-stationary, such that the post-intervention weight matrix $\bm{B}_{\bm{a}}$ varies over time. Although the proposed CSL-UCB algorithm keeps updating the causal structure, it should only utilize data generated by the current causal model. In this section, we develop a causal sub-graph change detector, which enables the proposed scheme to maintain a high cumulative reward even when the causal structure is non-stationary.
Consider a piece-wise stationary model, where the causal structure remains the same between two consecutive change-points. Without loss of generality, we focus on detecting the first change-point in time, denoted by $t_{c}$. To determine whether a change occurs at step $s$, a straightforward approach is to compare the distribution of observed values during $[1,s]$ with the distribution during $(s,t]$. If the distributions differ significantly, $s$ is claimed to be a change-point. However, the distribution shift of $\bm{x}$ only indicates the existence of a change, but not the source of the change. To understand this, recall the data generating process,
@@
where the distribution shift of $x_{i}$ could be due to the change of causal mechanism $\left[\bm{B}_{\bm{a}}\right]_{ki}$ or/and the distribution shifts of its parents $x_{k}$. As the source of change is not identifiable, the whole graph has to be re-learned.
Since causal mechanisms are modular and independent [25], the sparse mechanism shift hypothesis [34] suggests that changes in the data distribution generally are caused by changes in only a subset of causal mechanisms. In this case, re-learning the changed part is sufficient. To identify the set of changed mechanisms, we rely on the whitening transformation of $\bm{x}$,
$\bm{y}$ $\doteq(\bm{I}-\hat{\bm{B}}_{\bm{a}}^{t})^{T}\bm{x}=(\bm{I}-\bm{B}_{\bm{a}})^{T}\bm{x}-(\Delta\bm{B}_{\bm{a}}^{t})^{T}\bm{x}$ (39)
$=\bm{\epsilon}-(\Delta\bm{B}_{\bm{a}}^{t})^{T}\bm{x}.$ (40)
According to (40), $y_{i}$ and $\epsilon_{i}$ differ in distribution if and only if $\bigl{[}\Delta\bm{B}_{\bm{a}}^{t}\bigr{]}_{i}\neq\bm{0}$, in which case the estimated sub-graph model fails to explain the observed data well. For large discrepancy between distributions, there are mainly two possible reasons: causal structure change and FN errors in estimation. Thus apart from detecting causal structure changes, checking consistency between $\bm{y}$ and $\bm{\epsilon}$ also enables us to discover FN errors.
If the sub-graph associated with node $i$ changes at a time, the whitened observation $y_{i}^{\tau}$ would differ in distribution with $\epsilon_{i}$. Thus, to determine whether there is a change-point, we formulate a hypothesis testing problem,
@@
where $\mathcal{L}(\cdot;\nu,\sigma^{2})$ denotes the likelihood function of a Gaussian distribution with mean $\nu$ and variance $\sigma^{2}$. Note that $\Psi_{i}(s)$ essentially measures the gain in likelihood, by explaining the observed data during $(s,t]$ with an alternative distribution. Based on the statistic, the estimated change-point is given as
@@
$\quad=\sum_{\tau=s+1}^{t}\log\left[\frac{\sigma_{i}}{\hat{\sigma}_{i}}\cdot\exp\Bigl{(}\frac{(y_{i}^{\tau}-\nu_{i})^{2}}{2\sigma_{i}^{2}}-\frac{(y_{i}^{\tau}-\hat{\nu}_{i})^{2}}{2\hat{\sigma}_{i}^{2}}\Bigr{)}\right]$ (48)
@@
Note that $\Psi_{i}(s)$ is computed with samples in $(s,t]$, as shown in (47). As a result, there exists strong positive correlation the statistics. For example, $\Psi_{i}(s_{1})$ and $\Psi_{i}(s_{2})$ are computed with overlapped samples in the interval $(\max(s_{1},s_{2}),t]$. Therefore, we approximate the false alarm probability as
@@
While our proposed method may not be optimal, we emphasize that the approach is able to detect changes at the sub-graph level, resulting in improved sample and computational efficiency.
@@ VI Numerical Results
In this section, we numerically evaluate the performance of the proposed schemes in terms of both graph identification and intervention selection. Both stationary and non-stationary causal structures are considered for evaluation. For each Monte Carlo run, the causal structure is randomly generated, with the edge weights randomly sampled from the uniform distribution $\mathcal{U}(-2,2)$. The exogenous variables are independently sampled in each time step from the Gaussian distribution $\mathcal{N}(1,1)$. To estimate the empirical mutual information, the $k$-nearest neighbour distances based approach is employed [37]. For each set of parameters, we repeat the Monte Carlo run $M=100$ times.
@@ VI-A Performance for Causal Graph Identification
$\operatorname*{arg\,min}_{\hat{\bm{B}}\in\mathbb{R}^{N\times N}}\Bigl{\{}-\log\mathcal{L}(\bm{X}^{1:t};\hat{\bm{B}})+\beta_{1}\bigl{\lVert}\hat{\bm{B}}\bigr{\rVert}_{1}+\beta_{2}\bigl{[}\mathrm{tr}\bigl{(}e^{\hat{\bm{B}}\circ\hat{\bm{B}}}\bigr{)}-N\bigr{]}\Bigr{\}},$ (57)
where $\beta_{1}$, $\beta_{2}$ are the penalty coefficients, $\hat{\bm{B}}\circ\hat{\bm{B}}$ denotes the Hadamard product and the negative log-likelihood is
$-\log\mathcal{L}(\bm{X}^{1:t};\hat{\bm{B}})=\frac{N}{2}\log\biggl{(}\sum_{i=1}^{N}\sum_{\tau=1}^{t}\bigl{(}x_{i}^{\tau}-\bigl{[}\hat{\bm{B}}\bigr{]}_{i}^{T}\bm{x}^{\tau}\bigr{)}^{2}\biggr{)}$
$\hskip 140.0pt-\log\bigl{|}\mathrm{det}(\bm{I}-\hat{\bm{B}})\bigr{|}.$ (58)
@@
$\operatorname*{arg\,min}_{\hat{\bm{B}}\in\mathbb{R}^{N\times N}}\Bigl{\{}-\log\mathcal{L}(\bm{X}^{1:t};\hat{\bm{B}})+\beta_{1}\bigl{\lVert}\hat{\bm{B}}\bigr{\rVert}_{1}$
$\hskip 60.0pt-\log\mathrm{det}(\beta_{3}\bm{I}-\hat{\bm{B}}\circ\hat{\bm{B}})+N\log\beta_{3}\Bigr{\}},$ (59)
@@
Fig. 2 plots the graph FN rate as a function of the graph size, with two sample sizes, $T=200$ and $400$. The graph FN rate is defined as
$\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\bigl{[}\exists(i,j)\in\mathcal{V}\times\mathcal{V}:(i,j)\in\mathcal{B}(m),(i,j)\notin\hat{\mathcal{B}}(m)\bigr{]},$ (60)
where $\mathcal{B}(m)$, ^$\hat{\mathcal{B}}(m)$ represent the true and estimated edge sets in the $m$-th Monte Carlo run. Observe that the CSL-nw scheme has a graph FN rate of $100\%$ in all cases, which indicates that without edge weight regularization, the estimated graphs always have FN errors. The GOLEM scheme performs slightly better than the DAGMA scheme, while both of them perform much worse than the proposed CSL scheme. This is due to the fact that GOLEM and DAGMA do not distinguish between FP and FN errors, while the CSL scheme is designed to explicitly consider the FN rate. On average, the CSL scheme achieves a graph FN rate that is lower than the GOLEM scheme by $84.4\%$ and the DAGMA scheme by $87.9\%$. Interestingly, given a fixed number of samples, as the graph size increases, the FN rates for both GOLEM and DAGMA increase rapidly. In contrast, The CSL scheme does not suffer much from the increased graph size, as learning sub-graphs has higher sample efficiency.
@@
For completeness, we also include performance for both types of graph errors, measured by the normalized structure Hamming distance (SHD), defined as
$d(\mathcal{B},\hat{\mathcal{B}})\doteq\frac{1}{|\mathcal{B}|}\sum_{(i,j)\in\mathcal{V}\times\mathcal{V}}\Bigl{\{}\mathbb{I}\bigl{[}(i,j)\notin\mathcal{B},(i,j)\in\hat{\mathcal{B}}\bigr{]}+$
$\hskip 13.0pt\mathbb{I}\bigl{[}(i,j)\in\mathcal{B},(i,j)\notin\hat{\mathcal{B}}\bigr{]}-\mathbb{I}\bigl{[}(i,j)\in\mathcal{B},(j,i)\in\hat{\mathcal{B}}\bigr{]}\Bigr{\}}.$ (61)
In Fig. 3, the normalized SHD is plotted as a function of the graph size. We observe that in terms of normalized SHD, the CSL scheme does not perform significantly better than the GOLEM and DAGMA schemes. Particularly when the graph is small ($N=6$), DAGMA achieves the lowest normalized SHD for $200$ samples, while GOLEM achieves the lowest SHD for $400$ samples. Since both GOLEM and DAGMA learn the whole graph with gradient-based optimization, a sufficient amount of samples are necessary to achieve desirable performance. As the graph size increases, CSL offers strongly superior performance due to its improved sample complexity.

### v2

@@ IV Intervention Selection under Uncertainty
$\Delta\bm{B}_{\bm{a}}^{t}\doteq\widehat{\bm{B}}_{\bm{a}}^{t}-\bm{B}_{\bm{a}},\quad\Delta\bm{\mu}_{\bm{a}}^{t}\doteq\hat{\bm{\mu}}_{\bm{a}}^{t}-\bm{\mu}_{\bm{a}},$ (29)
where $\hat{\bm{\mu}}_{\bm{a}}^{t}$ represents the estimated mean of $\bm{x}$. As analyzed in Section III-C, the weight errors could have non-zero means, due to missing edges (false negatives). With the centered error matrix defined as
@@
$\left|\left[\Delta\bm{\mu}_{\bm{a}}^{t}\right]_{N}\right|$ $=\left|{\bigl{[}(\bm{I}-\widehat{\bm{B}}_{\bm{a}}^{t})^{-1}\bigr{]}}^{\top}_{N}{(\Delta\bm{B_{a}}^{t})}^{\top}\bm{\mu}_{\bm{a}}\right|$ (31)
$\leq\left|{\bigl{[}(\bm{I}-\widehat{\bm{B}}_{\bm{a}}^{t})^{-1}\bigr{]}}^{\top}_{N}{\mathbb{E}\bigl{[}\Delta\bm{B}_{\bm{a}}^{t}\big{|}\bm{X}^{1:t}(\bm{a})\bigr{]}}^{\top}\bm{\mu}_{\bm{a}}\right|$
$\hskip 40.0pt+\left|{\bigl{[}(\bm{I}-\widehat{\bm{B}}_{\bm{a}}^{t})^{-1}\bigr{]}}^{\top}_{N}{\bigl{(}\Delta\tilde{\bm{B}}_{\bm{a}}^{t}\bigr{)}}^{\top}\bm{\mu}_{\bm{a}}\right|.$ (32)
Recall that we provide an expression for the non-zero elements of $\mathbb{E}\bigl{[}\Delta\bm{B}_{\bm{a}}^{t}\big{|}\bm{X}^{1:t}(\bm{a})\bigr{]}$ in (28), which depends on the correlation between correctly declared parents and mistakenly rejected parents. However, since we are not able to know whether and where causal edges are rejected by mistake, the uncertainty bound cannot be computed at run-time.
@@
$\mathbb{P}\biggl{\{}\sigma_{\max}(\Delta\bm{B}_{\bm{a}}^{t})\geq 2\left(N^{2}+2N\right)^{1/4}\\ \times\sqrt{\ln\Bigl{(}\frac{2N}{\delta}\Bigr{)}\sum_{i=1}^{N}\lambda_{\max}(\bm{\Phi}_{\bm{a}}^{t}(i))}\biggr{\}}\leq\delta,$ (34)
@@
The proof relies on a key dilation transformation, Lemma 1 and the matrix Bernstein inequality [40]. Based on Lemma 2, the following theorem provides a concentration inequality for the error of the estimated reward.
@@
Under the assumption of MMSE estimation and no false negative errors, the following inequality holds:
@@
$U(\bm{X}^{1:t},\bm{a},\delta)=2\left(N^{2}+2N\right)^{1/4}\left\lVert\bigl{[}(\bm{I}-\widehat{\bm{B}}_{\bm{a}}^{t})^{-1}\bigr{]}_{N}\right\rVert_{2}\\ \times\left\lVert\bm{\mu}_{\bm{a}}\right\rVert_{2}\ \sqrt{\ln\Bigl{(}\frac{2N}{\delta}\Bigr{)}\sum_{i=1}^{N}\lambda_{\max}(\bm{\Phi}_{\bm{a}}^{t}(i))}.$ (36)
@@
The proof is based on the Woodbury matrix identity [41] and Lemma 2. Since the uncertainty of reward estimation cannot be measured in the presence of FN errors, we employ (35) as an approximation for the general case. The bound on uncertainty enables us to balance between exploration and exploitation, so that the cumulative regret is minimized in the long run. In the presence of FN errors, although the uncertainty bound may not hold exactly, the observed low cumulative regret in Section VI-B suggests, at least numerically, that the uncertainty bound has utility in our proposed algorithm.
@@
$\bm{a}^{t+1}=\operatorname*{arg\,max}_{\bm{a}}\left\{{\bigl{[}(\bm{I}-\widehat{\bm{B}}_{\bm{a}}^{t})^{-1}\bigr{]}}^{\top}_{N}\ \bm{\nu}+\alpha U(\bm{X}^{1:t},\bm{a},\delta)\right\},$ (37)
@@ V Causal Sub-graph Change Detection
So far, we have investigated the causal bandit problem, assuming that the underlying causal structure is invariant. However, it could be non-stationary in real-world scenarios, such that the post-intervention weight matrix $\bm{B}_{\bm{a}}$ varies over time. Although the proposed CSL-UCB algorithm keeps updating the causal structure, it should only utilize data generated by the current causal model. In this section, we develop a causal sub-graph change detector for both weight and topology changes, which enables the proposed scheme to maintain a high cumulative reward even when the causal structure is non-stationary.
Consider a piece-wise stationary model where the causal structure remains the same between two consecutive change points. Without loss of generality, we focus on detecting the first change-point in time, denoted by $t_{c}$. To determine whether a change occurs at step $s$, a straightforward approach compares the distribution of observed values during $[1,s]$ with the distribution during $(s,t]$. If the distributions differ significantly, $s$ is claimed to be a change point. However, the distribution shift of $\bm{x}$ only indicates the existence of a change but not the source of the change. To understand this, recall the data generating process,
@@
where the distribution shift of $x_{i}$ could be due to the change of causal mechanism $\left[\bm{B}_{\bm{a}}\right]_{i}$ or/and the distribution shifts of its parents $x_{k}$. As the source of change is not identifiable, the whole graph has to be re-learned.
Since causal mechanisms are modular and independent [31], the sparse mechanism shift hypothesis [42] suggests that changes in the data distribution generally are caused by changes in only a subset of causal mechanisms. In this case, it is sufficient to only re-learn the changed part. We rely on the whitening transformation of $\bm{x}$ to identify the set of changed mechanisms,
$\bm{y}$ $\doteq{(\bm{I}-\widehat{\bm{B}}_{\bm{a}}^{t})}^{\top}\bm{x}={(\bm{I}-\bm{B}_{\bm{a}})}^{\top}\bm{x}-{(\Delta\bm{B}_{\bm{a}}^{t})}^{\top}\bm{x}$ (39)
$=\bm{\epsilon}-{(\Delta\bm{B}_{\bm{a}}^{t})}^{\top}\bm{x}.$ (40)
According to (40), $y_{i}$ and $\epsilon_{i}$ differ in distribution if and only if $\bigl{[}\Delta\bm{B}_{\bm{a}}^{t}\bigr{]}_{i}\neq\bm{0}$, in which case the estimated sub-graph model fails to explain the observed data well. There are two main possible reasons for a large discrepancy between distributions: causal structure change and FN errors in estimation. Thus, besides detecting causal structure changes, checking the consistency between $\bm{y}$ and $\bm{\epsilon}$ also enables us to discover FN errors.
When the sub-graph associated with node $i$ changes, the whitened variable $y_{i}^{\tau}$ would differ in distribution from $\epsilon_{i}$. Thus, to determine whether there is a change point, we formulate a hypothesis testing problem,
@@
where $\mathcal{L}(\cdot;\nu,\sigma^{2})$ denotes the likelihood function of a Gaussian distribution with mean $\nu$ and variance $\sigma^{2}$. Note that $\Psi_{i}(s)$ essentially measures the gain in likelihood by explaining the observed data during $(s,t]$ with an alternative distribution. Based on the GLR statistic, the estimated change point is given as
@@
$\quad=\sum_{\tau=s+1}^{t}\log\left[\frac{\sigma_{i}}{\hat{\sigma}_{i}}\exp\Bigl{(}\frac{(y_{i}^{\tau}-\nu_{i})^{2}}{2\sigma_{i}^{2}}-\frac{(y_{i}^{\tau}-\hat{\nu}_{i})^{2}}{2\hat{\sigma}_{i}^{2}}\Bigr{)}\right]$ (48)
@@
Note that $\Psi_{i}(s)$ is computed with samples in $(s,t]$, as shown in (47). As a result, a strong positive correlation exists between the statistics calculated using overlapping subsets of the samples. For example, $\Psi_{i}(s_{1})$ and $\Psi_{i}(s_{2})$ are computed with overlapped samples in the interval $(\max(s_{1},s_{2}),t]$. Therefore, we approximate the false alarm probability as
@@
While our proposed method may not be optimal, we emphasize that the approach can detect changes at the sub-graph level, resulting in improved sample and computational efficiency.
@@ VI Numerical Results
In this section, we numerically evaluate the performance of the proposed schemes in terms of both graph identification and intervention selection. Both stationary and non-stationary causal structures are considered for evaluation. For each Monte Carlo run, the causal structure is randomly generated, with the edge weights randomly sampled from the uniform distribution $\mathcal{U}(-2,2)$. The exogenous variables are independently sampled in each time step from the Gaussian distribution $\mathcal{N}(1,1)$. To estimate the empirical mutual information, the $k$-nearest neighbor distances based approach is employed [45]. For each set of parameters, we repeat the Monte Carlo run $M=100$ times. [^2] The code is available at https://github.com/CalixPeng/Causal_Bandit.
@@ VI-A Performance for Causal Graph Identification
$\operatorname*{arg\,min}_{\widehat{\bm{B}}\in\mathbb{R}^{N\times N}}\Bigl{\{}-\log\mathcal{L}(\bm{X}^{1:t};\widehat{\bm{B}})+\beta_{1}\bigl{\lVert}\widehat{\bm{B}}\bigr{\rVert}_{1}+\beta_{2}\bigl{[}\mathrm{tr}\bigl{(}e^{\widehat{\bm{B}}\circ\widehat{\bm{B}}}\bigr{)}-N\bigr{]}\Bigr{\}},$ (57)
where $\beta_{1}$ and $\beta_{2}$ are the penalty coefficients, $\circ$ denotes the Hadamard product and the negative log-likelihood is
$-\log\mathcal{L}(\bm{X}^{1:t};\widehat{\bm{B}})=\frac{N}{2}\log\biggl{(}\sum_{i=1}^{N}\sum_{\tau=1}^{t}\bigl{(}x_{i}^{\tau}-{\widehat{\bm{B}}}^{\top}_{i}\bm{x}^{\tau}\bigr{)}^{2}\biggr{)}$
$\hskip 140.0pt-\log\bigl{|}\mathrm{det}(\bm{I}-\widehat{\bm{B}})\bigr{|}.$ (58)
@@
$\operatorname*{arg\,min}_{\widehat{\bm{B}}\in\mathbb{R}^{N\times N}}\Bigl{\{}-\log\mathcal{L}(\bm{X}^{1:t};\widehat{\bm{B}})+\beta_{1}\bigl{\lVert}\widehat{\bm{B}}\bigr{\rVert}_{1}$
$\hskip 60.0pt-\log\mathrm{det}(\beta_{3}\bm{I}-\widehat{\bm{B}}\circ\widehat{\bm{B}})+N\log\beta_{3}\Bigr{\}},$ (59)
@@
Fig. 2 plots the graph FN rate as a function of graph sizes, with two sample sizes, $T=200$ and $400$. The graph FN rate is defined as
$\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\bigl{(}\exists(i,j)\in\mathcal{V}\times\mathcal{V}:\bm{B}_{ij}(m)\neq 0,\widehat{\bm{B}}_{ij}(m)=0\bigr{)},$ (60)
where $\bm{B}(m)$ and $\widehat{\bm{B}}(m)$ represent the true and estimated edge weight matrices in the $m$-th Monte Carlo run. Observe that the CSL-nw scheme has a graph FN rate of $100\%$ in all cases, which indicates that without edge weight regularization, the estimated graphs always have FN errors. The GOLEM scheme performs slightly better than the DAGMA scheme, while both of them perform much worse than the proposed CSL scheme. This is due to the fact that GOLEM and DAGMA do not distinguish between FP and FN errors, while the CSL scheme is designed to explicitly consider the FN rate. On average, the CSL scheme achieves a graph FN rate that is lower than the GOLEM scheme by $81.1\%$ and the DAGMA scheme by $86.9\%$. Interestingly, given a fixed number of samples, as the graph size increases, the graph FN rates for both GOLEM and DAGMA increase rapidly. In contrast, The CSL scheme does not suffer much from the increased graph size, as learning sub-graphs has higher sample efficiency.
@@
To better understand the balance between false positives and false negatives, we consider precision and recall for evaluation, which are defined as
$\mathrm{precision}$ $\doteq\frac{\sum_{i,j}\mathbb{I}\bigl{(}\bm{B}_{ij}\neq 0,\widehat{\bm{B}}_{ij}\neq 0\bigr{)}}{\sum_{i,j}\mathbb{I}\bigl{(}\widehat{\bm{B}}_{ij}\neq 0\bigr{)}},$ (61)
$\mathrm{recall}$ $\doteq\frac{\sum_{i,j}\mathbb{I}\bigl{(}\bm{B}_{ij}\neq 0,\widehat{\bm{B}}_{ij}\neq 0\bigr{)}}{\sum_{i,j}\mathbb{I}\bigl{(}\bm{B}_{ij}\neq 0\bigr{)}}.$ (62)
Note that precision and recall are inversely correlated with the numbers of FP and FN, respectively. Figure 3 and 4 plot precision and recall as functions of graph sizes. We observe that the proposed CSL scheme achieves a close to one recall by focusing on the reduction of FN errors. On average, the recall rates of GOLEM, DAGMA and CSL are $92.7\%$, $93.8\%$ and $99.6\%$. On the other hand, the CSL scheme does not perform significantly better than GOLEM and DAGMA in terms of precision, especially for small graphs. The average precision rates achieved by GOLEM, DAGMA and CSL are $93.5\%$, $90.9\%$ and $96.7\%$. We also observe that both precision and recall of the CSL scheme increase as the graph becomes larger, while GOLEM and DAGMA have the opposite trend. It suggests that for the CSL scheme, as $N$ increases, the number of mistaken edges increases more slowly than the totally number of edges, which increases quadratically in $N$.
Figure 3: Precision as a function of graph sizes.
Figure 4: Recall as a function of graph sizes.
For completeness, we also include performance for both types of graph errors, measured by the normalized structural Hamming distance (NSHD), which is defined as
$\mathrm{NSHD}\doteq\sum_{i,j}\Bigl{[}$ $\mathbb{I}\bigl{(}\bm{B}_{ij}=0,\widehat{\bm{B}}_{ij}\neq 0\bigr{)}+\mathbb{I}\bigl{(}\bm{B}_{ij}\neq 0,\widehat{\bm{B}}_{ij}=0\bigr{)}$
$-\mathbb{I}\bigl{(}\bm{B}_{ij}\neq 0,\widehat{\bm{B}}_{ij}=0,\widehat{\bm{B}}_{ji}\neq 0\bigr{)}\Bigr{]}/N^{2}.$ (63)
In Fig. 5, the NSHD is plotted as a function of graph sizes. We observe that in terms of NSHD, the CSL scheme does not perform significantly better than the GOLEM and DAGMA schemes. Particularly when the graph is small ($N=6$), DAGMA achieves the lowest normalized SHD for $200$ samples, while GOLEM achieves the lowest SHD for $400$ samples. Since both GOLEM and DAGMA learn the whole graph with gradient-based optimization, a sufficient amount of samples are necessary to achieve desirable performance. As the graph size increases, CSL offers strongly superior performance due to its need for fewer samples than the other methods.